# Interpolation theorems for conjugations and applications

### v1

@@ Abstract.
Let $H$ be a separable complex Hilbert space. A conjugate-linear map $C:H\to H$ is a conjugation if it is an involutive isometry. In this paper, we consider the following interpolation problems: Let $\{x_{i}\}_{i\in I}$ and $\{y_{i}\}_{i\in I}$ be two orthogonal sets of vectors in $H$, and let $N$ and $\{N_{k}\}_{k\in K}$ be normal operators such that the $N_{k}$’s mutually commute. Then, under which conditions does there exist a conjugation $C$ on $H$ such that
@@
    $Cx_{i}=y_{i}$, for every $i\in I$, and $CNC=-N^{*}$.
We provide complete answers to problems (a) and (b). As a consequence of our results, we give necessary and sufficient conditions for the existence of solutions of some equations in $L^{\infty}(\mu)$.
@@ 1. Introduction
Throughout this paper, $H$ and $K$ are separable complex Hilbert spaces. We use the notation .,$\langle$.,$\rangle$ . , . ⟩ for the inner product of both $H$ and $K$. The algebra of all bounded linear operators on $H$ is denoted by ${\mathcal{B}}(H)$, and the identity operator on $H$ is denoted by $I_{H}$; if there is no ambiguity, we simply write $I$. The range and the null space of an operator $T$ are denoted by ${\rm ran}(T)$ and ${\rm ker}(T)$, respectively. All direct sums in this paper are orthogonal.
A conjugation $C$ on $H$ is a conjugate-linear map $C:H\to H$ that satisfies
@@
    $\langle Ch,Ck\rangle=\langle k,h\rangle$ for all $h,k\in H$.
The simplest example of a conjugation on $H$ is the map given by $C\left(\sum_{i}\lambda_{i}e_{i}\right)=\overline{\lambda_{i}}e_{i}$, where $\{e_{i}\}_{i}$ is an orthonormal basis of $H$. In fact, according to [13, Lemma 1], any conjugation can be represented in this manner with respect to some orthonormal basis.
@@
An operator $T\in{\mathcal{B}}(H)$ is called complex-symmetric (resp. skew-symmetric) if there exists a conjugation $C$ on $H$ such that $CTC=T^{*}$ (resp. $CTC=-T^{*}$); in this case, we say more precisely that $T$ is $C$-symmetric (resp. $C$-skew symmetric). An equivalent definition is that $T$ has a symmetric (resp. skew-symmetric) matrix $M$ with respect to some orthonormal basis; i.e., $M^{\rm tr}=M$ (resp. $M^{\rm tr}=-M$) where ${\rm tr}$ stands for the transpose ([14, 17]). These classes of linear operators have received significant attention in the last two decades. Applications of complex symmetric operators and skew-symmetric matrices are found in various areas (see, for instance, [13, 14, 15, 24] and [7, 22, 23]).
@@
In [18, 19, 26, 29], the authors investigated various interpolation problems for conjugations. In [18], Liu et al. established that for two orthogonal projections $P$ and $Q$ in ${\mathcal{B}}(H)$, a conjugation $C$ on $H$ exists with $CPC=Q$ if and only if ${\rm ran}(P)$ and ${\rm ran}(Q)$ are in symmetric position; i.e.,
$\dim\left({\rm ran}(P)^{\perp}\cap{\rm ran}(Q)\right)=\dim\left({\rm ran}(P)\cap{\rm ran}(Q)^{\perp}\right).$
In [26], Wang et al. proved that for two commuting self-adjoint operators $P,Q\in{\mathcal{B}}(H)$, there exists a conjugation $C$ on $H$ satisfying $CPC=Q$ if and only if there exists an isometric isomorphism $U$ between Hilbert spaces such that
$UPU^{-1}=B_{1}\oplus B_{2}\oplus A\quad\mbox{and}\quad UQU^{-1}=B_{2}\oplus B_{1}\oplus A$
where $B_{1}$, $B_{2}$ and $A$ are self-adjoint operators such that $B_{1}$ and $B_{2}$ have the same underlying Hilbert space and commute.
It is well-known that for every normal operator $N$, there exists a conjugation $C$ such that $CNC=N^{*}$ ([13]). However, this is not always the case if it is also required that $Cx_{i}=y_{i}$ for some families of vectors $\{x_{i}\}_{i\in I}$ and $\{y_{i}\}_{i\in I}$. It is the purpose of this paper to investigate such a problem. The paper is structured as follows.
In Section 2, we present our primary results regarding the interpolation problem: when does there exist a conjugation $C$ on $H$ satisfying $Cx_{i}=y_{i}$ and $CN_{k}C=N_{k}^{*}$ for all $i\in I$ and $k\in K$. Here, $\{x_{i}\}_{i\in I}$ and $\{y_{i}\}_{i\in I}$ are orthogonal subsets of $H$, and $\{N_{k}\}_{k\in K}$ is a collection of mutually commuting normal operators on $H$. We establish geometric conditions to answer this question, along with several equivalent conditions. As a consequence of these results, we extend a recent result by Mashreghi et al. [20] concerning hyperinvariant subspaces of unitary operators (see Proposition 2.5).
In Section 3, we investigate a “skew-symmetric” version of the earlier interpolation problem. We aim to identify the conditions under which there exists a conjugation $C$ on $H$ that can fulfill the conditions $Cx_{i}=y_{i}$ and $CNC=-N^{*}$ for all $i\in I$, with $N$ being a normal operator. The simplest answer is obtained when $I$ is finite.
In the last section, we apply our main results to give necessary and sufficient conditions for the existence of solutions of certain equations in $L^{\infty}(\mu)$ with coefficients in $L^{2}(\mu)$.
@@ 2. Interpolation theorems for conjugations : C-symmetric versions
Recall that a linear (or conjugate-linear) operator $V$ on $H$ is called partial isometry if the restriction of $V$ to ${\rm ker}(V)^{\perp}$ is an isometry.
@@ 2.1. Main results
Let $\mathcal{M}$ be an abelian von Neumann algebra of operators in ${\mathcal{B}}(H)$, and let $\{x_{i}\}_{i\in I}$ and $\{y_{i}\}_{i\in I}$ be two orthogonal sets of non-zero vectors in $H$. The following statements are equivalents:
@@
    There exists a conjugation $C$ on $H$ such that
    $CTC=T^{*}\;\;\mbox{and}\;\;Cx_{i}=y_{i}\;\;\;\mbox{for all $T\in\mathcal{M}$ and $i\in I$}.$
@@
    There exists a conjugate-linear partial isometry $V$ on $H$ such that
    $VT=T^{*}V,\;\;x_{i}\in{\rm ker}(V)^{\perp},\;\;Vx_{i}=y_{i}\;\;\mbox{and}\;\;Vy_{i}=x_{i}\;\;\;\mbox{for all $T\in\mathcal{M}$ and $i\in I$}.$
@@
    For every orthogonal projection $P\in\mathcal{M}$ and for all $i,j\in I$, we have
    $\langle Px_{i},x_{j}\rangle=\langle Py_{j},y_{i}\rangle\;\;\mbox{and}\;\;\langle Px_{i},y_{j}\rangle=\langle Px_{j},y_{i}\rangle.$
Remark 2.2.
Noting that $C(\lambda I)C=(\lambda I)^{*}$ for all conjugations $C$ and $\lambda\in\mathbb{C}$, one can see that when =$\mathcal{M}=\{\lambda I:\lambda\in\mathbb{C}\}$, we recapture [29, Theorem 2.1]. This theorem states that for two orthonormal sets $\{x_{i}\}_{i\in I}$ and $\{y_{i}\}_{i\in I}$ in $H$, there exists a conjugation $C$ on $H$ satisfying $Cx_{i}=y_{i}$ for every $i\in I$ if and only if $\langle x_{i},y_{j}\rangle=\langle x_{j},y_{i}\rangle$ for all $i,j\in I$. It should be noted that their result is important in proving our main results.
For a normal operator $N$, we denote by $E_{N}$ the associated spectral measure defined on Borel subsets of $\mathbb{C}$.
If $\mathcal{M}=W^{*}(N_{k}:k\in K)$, meaning that $\mathcal{M}$ is the smallest von Neumann algebra containing a family of mutually commuting normal operators $\{N_{k}\}_{k\in K}$, then the conditions outlined in Theorem 2.1 can be refined as follows.
Theorem 2.3.
Let $\{N_{k}\}_{k\in K}\subset{\mathcal{B}}(H)$ be a family of mutually commuting normal operators, and let $\{x_{i}\}_{i\in I}$ and $\{y_{i}\}_{i\in I}$ be two orthogonal sets of non-zero vectors in $H$. The following statements are equivalent:
  1. (i)
    There exists a conjugation $C$ on $H$ such that $CTC=T^{*}$ and $Cx_{i}=y_{i}$ for all $T\in W^{*}(N_{k}:k\in K)$ and $i\in I$.
  2. (ii)
    There exists a conjugation $C$ on $H$ such that $CN_{k}C=N_{k}^{*}$ and $Cx_{i}=y_{i}$ for all $k\in K$ and $i\in I$.
  3. (iii)
    There exists a conjugate-linear partial isometry $V$ on $H$ such that $VN_{k}=N_{k}^{*}V$, $x_{i}\in{\rm ker}(V)^{\perp}$, $Vx_{i}=y_{i}$ and $Vy_{i}=x_{i}$ for all $k\in K$ and $i\in I$.
  4. (iv)
@@
Given non-zero vectors $x,y\in H$, we denote by $x\otimes y$ the rank-one operator defined by $(x\otimes y)h=\langle h,y\rangle x$ for every $h\in H$.
If $I$ is of cardinal one, then Theorem 2.3 reduces to:
@@
Let $\{N_{k}\}_{k\in K}\subset{\mathcal{B}}(H)$ be a family of mutually commuting normal operators, and let $x,y\in H$ be non-zero vectors. The following statements are equivalent:
@@
    There exists a conjugation $C$ on $H$ satisfying $CN_{k}C=N_{k}^{*}$ and $Cx=y$ for every $k\in K$.
@@
    There exists a unitary operator $U\in{\mathcal{B}}(H)$ satisfying $UN_{k}=N_{k}U$ and $Ux=y$ for every $k\in K$.
@@
    There exists a partial isometry $V\in{\mathcal{B}}(H)$ satisfying $x\in{\rm ker}(V)^{\perp}$, $VN_{k}=N_{k}V$ and $Vx=y$ for every $k\in K$.
@@
    $\|x\|=\|y\|$ and there exists a conjugation $C$ on $H$ such that $N_{k}+\lambda x\otimes y$ is $C$-symmetric for all $k\in K$ and $\lambda\in\mathbb{C}$.
The proofs of Theorems 2.1, 2.3, and 2.4 will be given later in this section, following the necessary preliminary preparations.
A slight reformulation of [1, Theorem 2.1] shows that if $N\in{\mathcal{B}}(H)$ is a normal operator, $U\in{\mathcal{B}}(H)$ is a unitary operator commuting with $N$, and $x\in H$, then for every $\lambda\in\mathbb{C}$, there exists a conjugation $C$ on $H$ such that $N+\lambda x\otimes Ux$ is $C$-symmetric. It can be seen that the implication (ii)$\implies$(v) in Theorem 2.4 generalizes this result, as the conjugation $C$ is independent of $\lambda$.
A subspace $M$ of $H$ is said to be hyperinvariant for $T\in{\mathcal{B}}(H)$ if $M$ is invariant under every operator that commutes with $T$. Recently, in [20], Mashreghi et all. proved that $M$ is hyperinvariant for a unitary operator $U\in{\mathcal{B}}(H)$ if and only if $CM\subseteq M$ for every conjugation $C$ on $H$ satisfying $CUC=U^{*}$. As a consequence of Theorem 2.4, this result is extended to all normal operators.

### v3

@@ Abstract.
Let $\mathcal{H}$ be a separable complex Hilbert space. A conjugate-linear map $C:\mathcal{H}\to\mathcal{H}$ is called a conjugation if it is an involutive isometry. In this paper, we focus on the following interpolation problems: Let $\{x_{i}\}_{i\in I}$ and $\{y_{i}\}_{i\in I}$ be orthonormal sets of vectors in $\mathcal{H}$, and let $\{N_{k}\}_{k\in K}$ be a set of mutually commuting normal operators. We seek to determine under which conditions there exists a conjugation $C$ on $\mathcal{H}$ such that
@@
    $Cx_{i}=y_{i}$ and $CN_{k}C=-N_{k}^{*}$ for all $i\in I$ and $k\in K$.
We provide complete answers to problems (a) and (b) using the spectral projections of normal operators. Our results are then applied to the study of complex symmetric and skew symmetric operators, as well as to the characterization of hyperinvariant subspaces of normal operators through conjugations.
@@ 1. Introduction
Throughout this paper, $\mathcal{H}$ and $\mathcal{K}$ are separable complex Hilbert spaces. We use the notation .,$\langle$.,$\rangle$ . , . ⟩ for the inner product of both $\mathcal{H}$ and $\mathcal{K}$. The algebra of all bounded linear operators on $\mathcal{H}$ is denoted by ${\mathcal{B}}(\mathcal{H})$, and the identity operator on $\mathcal{H}$ is denoted by $I_{\mathcal{H}}$. The range and the null space of an operator $T$ are denoted by ${\rm ran}(T)$ and ${\rm ker}(T)$, respectively. All direct sums in this paper are orthogonal. The sets of non-negative integers and positive integers are denoted, as customary, by $\mathbb{N}$ and $\mathbb{Z}^{+}$, respectively. Throughout the paper, $\mathcal{N}=\{N_{k}\}_{k\in K}$ and =$\mathcal{M}=\{M_{k}\}_{k\in K}$ are two sets of normal operators on $\mathcal{H}$ such that each two operators in their union are commuting, and $X=\{x_{i}\}_{i\in I}$ and $Y=\{y_{i}\}_{i\in I}$ will always denote two orthonormal sets of non-zero vectors in $\mathcal{H}$.
Definition 1.1.
A conjugation $C$ on $\mathcal{H}$ is a conjugate-linear map $C:\mathcal{H}\to\mathcal{H}$ that satisfies
@@
    $\langle Ch,Ck\rangle=\langle k,h\rangle$ for all $h,k\in\mathcal{H}$.
The simplest example of a conjugation on $\mathcal{H}$ is the map given by $C\left(\sum_{i}\lambda_{i}e_{i}\right)=\overline{\lambda_{i}}e_{i}$ where $\{e_{i}\}_{i}$ is an orthonormal basis of $\mathcal{H}$. In fact, according to [14, Lemma 1], any conjugation can be represented in this manner with respect to some orthonormal basis.
@@
An operator $T\in{\mathcal{B}}(\mathcal{H})$ is called complex-symmetric (resp. skew-symmetric) if there exists a conjugation $C$ on $\mathcal{H}$ such that $CTC=T^{*}$ (resp. $CTC=-T^{*}$); in this case, we say more precisely that $T$ is $C$-symmetric (resp. $C$-skew symmetric). An equivalent definition is that $T$ has a symmetric (resp. skew-symmetric) matrix $M$ with respect to some orthonormal basis; i.e., $M^{\rm tr}=M$ (resp. $M^{\rm tr}=-M$) where ${\rm tr}$ stands for the transpose ([15, 18]). Denote by ${\rm CSO}$ and ${\rm SSO}$ the classes of complex symmetric and skew symmetric operators on $\mathcal{H}$, respectively. These classes of linear operators have received significant attention in the last two decades. Applications of complex symmetric operators and skew-symmetric matrices are found in various areas (see, for instance, [14, 15, 16, 25] and [8, 23, 24]).
Given two vectors $x,y\in\mathcal{H}$, we denote by $x\otimes y$ the operator defined by $(x\otimes y)h=\langle h,y\rangle x$ for every $h\in\mathcal{H}$. It is easy to see that if $C$ is a conjugation, then
(1.1) $C(x\otimes y)C=(Cx)\otimes(Cy)\quad\mbox{for all }x,y\in\mathcal{H}.$
@@
In [19, 20, 27, 30], the authors investigated various interpolation problems for conjugations. For instance, in [19], Liu et al. established that for two orthogonal projections $P$ and $Q$ in ${\mathcal{B}}(\mathcal{H})$, a conjugation $C$ on $\mathcal{H}$ exists with $CPC=Q$ if and only if ${\rm ran}(P)$ and ${\rm ran}(Q)$ are in symmetric position; i.e.,
$\dim\big{(}{\rm ran}(P)^{\perp}\cap{\rm ran}(Q)\big{)}=\dim\big{(}{\rm ran}(P)\cap{\rm ran}(Q)^{\perp}\big{)}.$
The main motivation for our results in Section 2 stems from the study of complex symmetric operators. It should be mentioned that the class ${\rm CSO}$ is not stable under the addition. Indeed, it is well-known that every normal operator $N\in{\mathcal{B}}(\mathcal{H})$ is complex symmetric. However, even its sum with the simplest complex symmetric operator may fail to be in ${\rm CSO}$ (see [1, Theorem 2.4]). The main results in Section 2 can be used to study complex symmetry of $N+T$ for a dense class of operators $T$ in ${\mathcal{B}}(\mathcal{H})$. To elaborate, Zhu and Li proved in [30] that operators $T\in{\mathcal{B}}(\mathcal{H})$ of the form
(1.2) $T=\sum_{j\in J}r_{j}e_{j}\otimes f_{j},$
where $\{r_{j}\}_{j\in J}$ are distinct positive scalars and $\{e_{j}\}_{j\in J}$ and $\{f_{j}\}_{j\in J}$ are orthonormal subsets of $\mathcal{H}$, form a dense class in ${\mathcal{B}}(\mathcal{H})$. Moreover, they showed that such an operator is complex symmetric if and only if there exist unimodular scalars $\lambda_{j}\in\mathbb{C}$, $j\in J$, such that
(1.3) $\lambda_{i}\langle e_{i},f_{j}\rangle=\lambda_{j}\langle e_{j},f_{i}\rangle\quad\mbox{for all }i,j\in J.$
Corollary 2.7 provides geometric conditions for the existence of a conjugation $C$ such that $N_{k}+\lambda T$ is $C$-symmetric for all $k\in K$ and $\lambda\in\mathbb{C}$. When $N_{k}=0$ for every $k\in K$, we recover the result in (1.3) with ease. Furthermore, our results extend a recent finding by Mashreghi et al. [21] on hyperinvariant subspaces of unitary operators (see Proposition 2.10).
In Section 3, we establish necessary and sufficient conditions for the existence of a conjugation $C$ such that $CN_{k}C=-N_{k}^{*}$ and $Cx_{i}=y_{i}$ for all $k\in K$ and $i\in I$. These results are used to characterize skew symmetry of operators of the form
(1.4) $N+\sum_{j\in J}r_{j}\left(e_{j}\otimes e_{j}-f_{j}\otimes f_{j}\right),$
where $N\in{\mathcal{B}}(\mathcal{H})$ is normal, $\{r_{j}\}_{j\in J}$ are distinct positive scalars and $\{e_{j},f_{j}\}_{j\in J}$ is an orthonormal subset. Notably, skew symmetric operators of the form (1.4) are dense in ${\rm SSO}$ (see Section 3).
To prove the main results of Section 3, we will uncover certain facts about skew symmetric normal operators. We will show that $\{N_{k}\}_{k\in K}$ are simultaneously skew-symmetric (i.e., there exists a conjugation $C$ such that $CN_{k}C=-N_{k}^{*}$ for every $k\in K$) if and only if there exists a unitary operator $U\in{\mathcal{B}}(\mathcal{H})$ such that $UN_{k}U^{*}=-N_{k}$ for every $k\in K$. Furthermore, for a normal operator $N\in{\mathcal{B}}(\mathcal{H})$, we provide a condition on the multiplicity of $N$ under which skew symmetry of $N\oplus T$ implies that one of $T$, where $T$ is an arbitrary operator.
@@ 2. Interpolation theorems for conjugations : $C$-symmetric versions
Given a subset $\mathcal{A}=\{A_{k}\}_{k\in K}\subset{\mathcal{B}}(\mathcal{H})$ of mutually commuting normal operators and a subset of vectors $Z\subseteq\mathcal{H}$, we denote
$L_{\mathcal{A},Z}=\bigvee\big{\{}\prod_{k\in K_{0}}A_{k}^{n_{k}}A_{k}^{m_{k}*}h:K_{0}\subseteq K\mbox{ is finite, }n_{k},m_{k}\in\mathbb{N}\mbox{ and }h\in Z\big{\}}$
where $\vee$ denotes the closed linear span. By the Fuglede Theorem, we also have $A_{k}A_{l}^{*}=A_{l}^{*}A_{k}$ for any $k,l\in K$. Hence, one can see that $L_{\mathcal{A},Z}$ is a reducing subspace of each $A_{k}$.
For a normal operator $N$, we denote by $E_{N}$ the associated spectral measure defined on Borel subsets of $\mathbb{C}$. The following theorem is central to proving our main results.
@@ 2.1. Main results
Assume that $\mathcal{H}=L_{\mathcal{N}\cup\mathcal{M},X\cup Y}$. The following statements are equivalent:
@@
    There exists a conjugation $C$ on $\mathcal{H}$ such that $CN_{k}C=M_{k}^{*}$ and $Cx_{i}=y_{i}$ for all $k\in K$ and $i\in I$.
@@
    For all finite subsets $K_{0},K_{1}\subseteq K$, Borel subsets $\{\Delta_{k}\}_{k\in K_{0}}$ and $\{D_{k}\}_{k\in K_{1}}$ of $\mathbb{C}$, and $i,j\in I$, we have
    $\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},\left(\prod_{l\in K_{1}}E_{M_{l}}(D_{l})\right)x_{j}\right\rangle=\left\langle\left(\prod_{l\in K_{1}}E_{N_{l}}(D_{l})\right)y_{j},\left(\prod_{k\in K_{0}}E_{M_{k}}(\Delta_{k})\right)y_{i}\right\rangle$
    and
    $\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},\left(\prod_{l\in K_{1}}E_{M_{l}}(D_{l})\right)y_{j}\right\rangle=\left\langle\left(\prod_{l\in K_{1}}E_{N_{l}}(D_{l})\right)x_{j},\left(\prod_{k\in K_{0}}E_{M_{k}}(\Delta_{k})\right)y_{i}\right\rangle.$
Remark 2.2.
Recall that an anti-automorphism $\rho:{\mathcal{B}}(\mathcal{H})\to{\mathcal{B}}(\mathcal{H})$ is a vector space automorphism satisfying $\rho(T^{*})=\rho(T)^{*}$ and $\rho(TS)=\rho(S)\rho(T)$ for all $T,S\in{\mathcal{B}}(\mathcal{H})$. It is well-known that each anti-automorphism $\rho:{\mathcal{B}}(\mathcal{H})\to{\mathcal{B}}(\mathcal{H})$, of order two (i.e., $\rho^{2}=I_{{\mathcal{B}}(\mathcal{H})}$), has the form $\rho(T)=CT^{*}C$ for every $T\in{\mathcal{B}}(\mathcal{H})$, where $C$ is a conjugation on $\mathcal{H}$. In the light of this fact, it is not hard to see that Theorem 2.1 can be used to obtain conditions under which there exists an anti-automorphism $\rho:{\mathcal{B}}(\mathcal{H})\to{\mathcal{B}}(\mathcal{H})$ of order two such that
$\rho(T)=T\quad\mbox{and}\quad\rho(N_{k})=M_{k}\;\mbox{for every }k\in K,$
where $T$ is an operator of the form (1.2).
For a subspace $M$ of $\mathcal{H}$, denote by $P_{M}\in{\mathcal{B}}(\mathcal{H})$ the orthogonal projection of $\mathcal{H}$ onto $M$. The following corollary is a direct consequence of Theorem 2.1.
Corollary 2.3.
Assume that =$\mathcal{H}=\vee\{x_{i},y_{i}:i\in I\}$, and let $M$ and $N$ be subspaces of $\mathcal{H}$ such that $P_{M}$ and $P_{N}$ commute. The following are equivalent:
  1. (i)
    There exists a conjugation $C$ on $\mathcal{H}$ such that $CM=N$ and $Cx_{i}=y_{i}$ for every $i\in I$.
  2. (ii)
    For all $i,j\in I$ and $r,s\in\{0,1\}$, we have
    $\left\langle P_{N}^{r}P_{M}^{s}x_{i},x_{j}\right\rangle=\left\langle P_{N}^{s}P_{M}^{r}y_{j},y_{i}\right\rangle\quad\mbox{and}\quad\left\langle P_{N}^{r}P_{M}^{s}x_{i},y_{j}\right\rangle=\left\langle P_{N}^{s}P_{M}^{r}x_{j},y_{i}\right\rangle.$
Recall that a linear (or conjugate-linear) operator $V$ on $\mathcal{H}$ is called partial isometry if the restriction of $V$ to ${\rm ker}(V)^{\perp}$ is an isometry.
If $N_{k}=M_{k}$ for every $k\in K$, Theorem 2.1 becomes:
Theorem 2.4.
The following statements are equivalent:
  1. (i)
    There exists a conjugation $C$ on $\mathcal{H}$ such that $CN_{k}C=N_{k}^{*}$ and $Cx_{i}=y_{i}$ for all $k\in K$ and $i\in I$.
  2. (ii)
    There exists a conjugate-linear partial isometry $V$ on $\mathcal{H}$ such that $VN_{k}=N_{k}^{*}V$, $x_{i}\in{\rm ker}(V)^{\perp}$, $Vx_{i}=y_{i}$ and $Vy_{i}=x_{i}$ for all $k\in K$ and $i\in I$.
@@
Remark 2.5.
Taking $N_{k}=0$, for every $k\in K$, in Theorem 2.4, we recapture [30, Theorem 2.1] which states that there exists a conjugation $C$ on $\mathcal{H}$ satisfying $Cx_{i}=y_{i}$ for every $i\in I$ if and only if $\langle x_{i},y_{j}\rangle=\langle x_{j},y_{i}\rangle$ for all $i,j\in I$.
Remark 2.6.
One can use Theorem 2.4 to show that if the subspaces $M$ and $N$ in Corollary 2.3 are equal, then the condition “=$\mathcal{H}=\vee\{x_{i},y_{i}:i\in I\}$” is no longer needed.
Corollary 2.7.
Let $\{r_{i}\}_{i\in I}$ be a set of distinct positive scalars, and let $T=\sum_{i\in I}r_{i}x_{i}\otimes y_{i}$. The following statements are equivalent:
  1. (i)
    There exists a conjugation $C$ on $\mathcal{H}$ such that $N_{k}+\lambda T$ is $C$-symmetric for all $k\in K$ and $\lambda\in\mathbb{C}$.
  2. (ii)
    There exist unimodular scalars $\lambda_{i}\in\mathbb{C}$, for $i\in I$, such that
    $\lambda_{i}\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},x_{j}\right\rangle=\lambda_{j}\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)y_{j},y_{i}\right\rangle$
    and
    $\lambda_{i}\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},y_{j}\right\rangle=\lambda_{j}\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{j},y_{i}\right\rangle$
    for all finite subsets $K_{0}\subseteq K$, Borel subsets $\{\Delta_{k}\}_{k\in K_{0}}$ of $\mathbb{C}$, and $i,j\in I$.
Proof.
(ii)$\implies$(i). For every $i\in I$, put $\widetilde{x_{i}}=\lambda_{i}x_{i}$. It is easy to see that the orthonormal sets $\{\widetilde{x_{i}}\}_{i\in I}$ and $\{y_{i}\}_{i\in I}$ satisfy the assertion (iii) in Theorem 2.4. Thus, there exists a conjugation $C$ on $\mathcal{H}$ such that $CN_{k}C=N_{k}^{*}$ and $C\widetilde{x_{i}}=y_{i}$ for all $k\in K$ and $i\in I$. Moreover, since $C^{-1}=C$, we also have $Cy_{i}=\widetilde{x_{i}}$. Now, by (1.1), one can check that $C(x_{i}\otimes y_{i})C=y_{i}\otimes x_{i}$ for every $i\in I$. Therefore, $CTC=T^{*}$, and consequently, $C(N_{k}+\lambda T)C=(N_{k}+\lambda T)^{*}$ for all $\lambda\in\mathbb{C}$ and $k\in K$.
(i)$\implies$(ii). Since the class of $C$-symmetric operators is a vector space and $\lambda$ is arbitrary, we have $CN_{k}C=N_{k}^{*}$ for every $k\in K$ and $CTC=T^{*}$. The later equality yields $CT^{*}TC=(CT^{*}C)(CTC)=TT^{*}$, and consequently $C(T^{*}T-r_{i}I_{\mathcal{H}})C=TT^{*}-r_{i}I_{\mathcal{H}}$ for every $i\in I$. In particular, $C{\rm ker}(TT^{*}-r_{i}I_{\mathcal{H}})\subseteq{\rm ker}(T^{*}T-r_{i}I_{\mathcal{H}})$ for every $i\in I$. Therefore, $C\left(\vee\{x_{i}\}\right)=\vee\{y_{i}\}$ for every $i\in I$. Since $C$ is isometric, there exist unimodular scalars $\lambda_{i}$, for $i\in I$, such that $C(\lambda_{i}x_{i})=y_{i}$ for every $i\in I$. Now, using Theorem 2.4, one can easily see that the scalars $\lambda_{i}$, $i\in I$, satisfy all the desired equalities. ∎
If $I$ is of cardinal one, then we get a linear version of Theorem 2.4 as follows:
@@
Let $x,y\in\mathcal{H}$ be unit vectors. The following statements are equivalent:
@@
    There exists a conjugation $C$ on $\mathcal{H}$ satisfying $CN_{k}C=N_{k}^{*}$ and $Cx=y$ for every $k\in K$.
@@
    There exists a unitary operator $U\in{\mathcal{B}}(\mathcal{H})$ satisfying $UN_{k}=N_{k}U$ and $Ux=y$ for every $k\in K$.
@@
    There exists a partial isometry $V\in{\mathcal{B}}(\mathcal{H})$ satisfying $x\in{\rm ker}(V)^{\perp}$, $VN_{k}=N_{k}V$ and $Vx=y$ for every $k\in K$.
@@
    There exists a conjugation $C$ on $\mathcal{H}$ such that $N_{k}+\lambda x\otimes y$ is $C$-symmetric for all $k\in K$ and $\lambda\in\mathbb{C}$.
The proofs of Theorems 2.1, 2.4, and 2.8 will be given later in this section, following the necessary preliminary preparations.
Remark 2.9.
A slight reformulation of [1, Theorem 2.1] shows that if $N\in{\mathcal{B}}(\mathcal{H})$ is a normal operator, $U\in{\mathcal{B}}(\mathcal{H})$ is a unitary operator commuting with $N$, and $x\in\mathcal{H}$, then for every $\lambda\in\mathbb{C}$, there exists a conjugation $C$ on $\mathcal{H}$ such that $N+\lambda x\otimes Ux$ is $C$-symmetric. It can be seen that the implication (ii)$\implies$(v) in Theorem 2.8 generalizes this result, as the conjugation $C$ is independent of $\lambda$.
A subspace $M$ of $\mathcal{H}$ is said to be hyperinvariant for $T\in{\mathcal{B}}(\mathcal{H})$ if $M$ is invariant under every operator that commutes with $T$. Recently, in [21], Mashreghi et all. proved that $M$ is hyperinvariant for a unitary operator $U\in{\mathcal{B}}(\mathcal{H})$ if and only if $CM\subseteq M$ for every conjugation $C$ on $\mathcal{H}$ satisfying $CUC=U^{*}$. As a consequence of Theorem 2.8, this result is extended to all normal operators.